OpenStudy (anonymous):

What is the equation of the line that passes through the points (-1, 3) and (1, 11)? A. 4x - y = 1 B. 4x + y = 1 C. x - 4y = -13 D. 4x - y = -7

OpenStudy (anonymous):

Did you do the math?

OpenStudy (tkhunny):

You could simply substitute the two points and see if they work. A. 4x - y = 1 (-1, 3) ==> 4(-1) - (3) = -4 - 3 = -7 and that is NOT 1 -- Forget about A B. 4x + y = 1 Demonstrate!

OpenStudy (anonymous):

http://www.hackmath.net/en/calculator/line-slope?x0=-1&y0=3&x1=1&y1=11&submit=Solve thus D is good answer: y = 4x+7

OpenStudy (anonymous):

Thank you :)